OpenStudy (nosnip):

Solving Algebraic Fractions 3/x + 3/2x = 2

OpenStudy (anonymous):

do you want to solve this equation \[\frac{ 3 }{ x }+\frac{ 3 }{ 2x }=2\]

OpenStudy (nosnip):

yes please

OpenStudy (anonymous):

\[\frac{ 3 }{ x }=\frac{ 6 }{ 2x }\]\[\frac{ 6 }{ 2x }+\frac{ 3 }{ 2x }=2\]\[\frac{ 9 }{ 2x }=2\]4x=9 \[x=\frac{ 9 }{ 4 }\]

OpenStudy (nosnip):

you don't need to times the =2 by 2x?

OpenStudy (nosnip):

when the other side is being multiplied by 2x

OpenStudy (nosnip):

Thank you!!